OpenStudy (anonymous):

pt. on graph of f(x)=x^2-1 closest to (-4,0)? i'm getting (-√3/2,1/2) but not sure

OpenStudy (anonymous):

That doesn't satisfy the f(x) so it can't be.

OpenStudy (anonymous):

Unless you mean sqrt(3/2) that is.

OpenStudy (anonymous):

yes there should be a sqrt symbol there? not visible?

OpenStudy (anonymous):

hey simon i need ur help

OpenStudy (anonymous):

with what?

OpenStudy (anonymous):

you missed brackets to make it what you wanted. I think you answer is wrong as it doesn't satisfy the optimisation equation.

OpenStudy (anonymous):

after ur done with this problem. could u help me with the problem that u answered for me. i whata kno how u got the answer step by step :) pretty plez

OpenStudy (anonymous):

i was minimizing 4x^3-2x+8

OpenStudy (anonymous):

flyqueen, maybe it was another simon?

OpenStudy (anonymous):

You shouldn't be minimising that, you should be finding the when it equals 0.

OpenStudy (anonymous):

oh lol ur rite ur the wrong simon sorry

OpenStudy (anonymous):

thats what i meant, and got x=-sqrt(3/2), y=1/2?

OpenStudy (anonymous):

Okay, now put -sqrt(3/2) into that equation. It doesn't equal 0.

OpenStudy (anonymous):

Answer is x = -1.3917... whatever that is in a nice form if there is one.

OpenStudy (anonymous):

how did you solve for zero? im getting 4x^3-2x+8=0 4x^3-2x=-8 2x(x^2-1)=-8 x=-4 or x=+-sqrt(-7)

OpenStudy (anonymous):

I put it into a calculator. If you cant spot a factorisation, cubics are typically hard to solve by hand. Use a calculator or the standard formula. I don't understand how you got to your answer given the line above it, but none of those x solve the equation.

OpenStudy (anonymous):

thanks